The Badgers are a pro sports team. The teams sells three different type of tickets: Deluxe, Lower level and upper level. Information from the prior year is as follows: (TIP- Revenue Variances)

BUDGET:

Contribution Margin

      Number of

          Per seat

           Seats

deluxe

$          100

           2,000

lower level

               50

           6,000

upper level

               20

           8,000

ACTUAL:

Contribution Margin

      Number of

          Per seat

           Seats

deluxe

$          110

           2,200

lower level

               50

           6,200

upper level

               20

           8,400

Total Market Size

Budget

        40,000

Actual

        50,000

Required:

  1. Calculate the Sales Volume Variance for each product.
  2. Calculate the Sales Quantity Variance for each product.
  3. Calculate the Market Share Variance in terms of contribution margin.
  4. Calculate the Market Size Variance in terms of contribution margin.
Answer #1

a. Sales Volume Variance"

Sales Volume Variance =(Actual Units-Budgeted units)*Budgeted Contribution per unit

Delux : (2200-2000)*$100=$20,000.... Favorable

Lower Level : (6200-6000)*$50=$10,000.....Favorable

Upper Level : (8400-8000)*$20=$8,000.....Favorable

b Sales Quantity Variance:

Sales Quantity Variance :

Total Budgeted Sales( units)=2000+6000+8000=16000

Proportion of Delux =2000/16000=0.125

Proportion of Lower Level =6000/16000=0.375

Proportion of Upper Level=8000/16000=0.5

Total Actual sales=2200+6200+8400=16800

UNIT SALES AT STANDARD MIX:

Delux : 0.125*16800=2100

Lower Level : 0.375*16800=6300

Upper Level : 0.5*16800=8400

Sales Quantity Variance =(Budgeted SalesQuantity -Unit sales Quantity at standard Mix)* Budgeted Unit Contribution

Delux : (2000-2100)*$100=$10,000 Favorable

Lower Level : (6000-6300)*$50=$15,000..Favorable

Upper Level : (8000-8400)$20=$8,000.Favorable

Market Size Variance :

Budgeted Sales =16000 units

Budgeted Market Size=40000

Budgeted market share =16000/40000=0.4

Budgeted unit Contribution =(2000*$100+6000*$50+8000*$20)/16000=$41.25

Market Size Variance =(Actual Market Size-Budgeted Market Size)*Market Share*Budgeted unit contribution

Market Size Variance =(50000-40000)*0.4*41.25=$165,000 Favorable

Market Share Variance =(Actual Market Share-Budgeted Market Share)*Actual Market Size* Budgeted unit contribution

Actual Sales =2200+6200+8400=16800

Actual Market size =50000

Actual Market Share =16800/50000=0.336

Market Share Variance =(0.336-0.4)*50000*$41.25=$132,000 Unfavorable